# Local wind speed forecasting at short time horizons relying on both Numerical Weather Prediction and observations from surrounding stations

### v1

@@ 5 Summary and prospects
We find that our hybrid model significantly outperforms baseline models, including raw NWP predictions (AROME and ARPEGE), persistence models, and linear regression models, across all forecast horizons (1-6 hours). When analyzing feature importance, we observe that wind components from NWP models are the most important features for prediction accuracy. AROME features remain consistently important across all horizons, while ARPEGE features gain importance at longer horizons. Direct observations from ground stations are more useful for forecasts at shortest horizons. We also show the probabilistic version of the model slightly outperforms the deterministic version in terms of RMSE. Fine-tuning the model for specific stations in Corsica further improves forecasting accuracy, particularly for smallest time horizons. This approach that leverages the optimal weights obtained from the global training phase mitigates overfitting when one wants to adapt the model to a specific location.
@@
In the context of extreme wind prediction, defined as the occurrence of wind speeds exceeding high quantiles (e.g., 90 %, 95%, 99%) of site-specific distributions, we demonstrate that the probabilistic model significantly outperforms linear regression models and post-processed AROME predictions, especially at longest forecasting horizons. We also show that our approach yields much more accurate predictions compared to its deterministic counterpart, which is based solely on the expected mean value. Moreover, the model’s performance improves as the quantile increases. This observation will have to be further investigated using a larger dataset, as it suggests that, in terms of PSS score, conditions leading to very strong wind events may be easier to anticipate than those associated with milder wind speeds.
@@
In conclusion, the hybrid neural network model demonstrates significant improvements in short-term wind speed forecasting by integrating NWP data with ground station observations. Its ability to provide both deterministic and probabilistic forecasts makes it a valuable tool for applications in renewable energy, public safety, and operational decision-making. A perspective for future work is focusing on the probabilistic aspects of the model to improve its reliability in terms of probabilistic metrics as logScore or CRPS (Baggio and Muzy, 2024). Our parametric approach can be directly compared with the non-parametric model based on normalizing flows proposed by (Marcille et al., 2024). Additionally, our predictions could be improved by leveraging, as pioneered by Rasp and Lerch (2018), ensemble NWP forecasting. This can potentially leads to better risk assessment for extreme events (Primo et al., 2024). Another promising direction involves incorporating additional data sources such as orography or surface roughness, that have been demonstrated to influence the reliability of wind speed predictions (Rasp and Lerch, 2018; Veldkamp et al., 2021). Finally, advancements in deep learning architectures offer appealing opportunities for further refinement. While LSTMs and CNNs have demonstrated strong performance, recent advances in Transformer-based models have shown great potential for capturing long-range dependencies in spatiotemporal data (see, e.g., Ødegaard Bentsen et al. (2023)).
Acknowledgements.
[…]
@@ Acknowledgements.
where the parameter $\lambda^{2}$ is referred to, in the literature on turbulence, as the “intermittency coefficient” Frisch (1995). As advocated in Baggio and Muzy (2024), this last formula is evaluated using the following Gauss-Hermite quadrature:
[…]
@@
with $n=11$ that is sufficient for the purpose of this paper (Baggio and Muzy, 2024).
Since for a Rice law of parameter $\nu$ and $\sigma^{2}$, the mean value is $\mu_{1}={\displaystyle\sigma{\sqrt{\frac{\pi}{2}}}\,\,L_{\frac{1}{2}}\left(-\frac{\nu^{2}}{2\sigma^{2}}\right)}$, where $L_{\frac{1}{2}}$ stands for the order $\frac{1}{2}$ Laguerre polynomial, the mean value of a M-Rice distribution reads:
[…]
$\simeq$ $\frac{\sigma}{\sqrt{2}}\sum_{i=1}^{n}w_{i}\,L_{\frac{1}{2}}\left(-\frac{e^{2\sqrt{2}\lambda y_{i}}\nu^{2}}{2\sigma^{2}}\right)$

### v2

@@ 7 Summary and prospects
The hybrid model proposed in this work significantly outperforms baseline models, including raw NWP predictions (AROME and ARPEGE), persistence models, and linear regression models, across all forecast horizons (1-6 hours). It also performs better than a neural network specifically optimized to handle recent observations at a given ground station and its neighboring sites as proposed in Baïle and Muzy (2023). When analyzing feature importance, it can be observed that wind components are the most important features for prediction accuracy. AROME features remain consistently important across all horizons, while ARPEGE features gain importance at longer horizons. Direct observations from ground stations are more useful for forecasts at shortest horizons. The probabilistic version of the model appears to slightly outperforms the deterministic version in terms of RMSE. Fine-tuning the model for specific stations further improves forecasting accuracy, particularly for smallest time horizons. This approach that leverages the optimal weights obtained from the global training phase mitigates overfitting when one wants to adapt the model to a specific location.
@@
In the context of extreme wind prediction, defined as the occurrence of wind speeds exceeding high quantiles (e.g., 90 %, 95%, 99%) of site-specific distributions, our results demonstrate that the probabilistic model significantly outperforms linear regression models and post-processed AROME predictions, especially at longer forecasting horizons. It is also shown that our approach yields much more accurate predictions compared to its deterministic counterpart, which is based solely on the expected mean value. Moreover, the model’s performance improves as the quantile increases. This observation will have to be further investigated using a larger dataset, as it suggests that, in terms of PSS score, conditions leading to very strong wind events may be easier to anticipate than those associated with milder wind speeds.
@@
In conclusion, the hybrid neural network model demonstrates significant improvements in short-term wind speed forecasting by integrating NWP data with ground station observations. Its ability to provide both deterministic and probabilistic forecasts makes it a valuable tool for applications in renewable energy, public safety, and operational decision-making. Many avenues can be explored to build upon this work. First, to better understand the proposed model limitations and guide targeted improvements, in continuation of the preliminary findings of Secs 5.1.2 and 5.2.2, a promising direction for future work is to investigate the specific conditions that influence the magnitude of prediction errors. Such an analysis could include evaluating conditional errors stratified by various key factors such as wind regime, terrain complexity, or seasonal variability as first experimented in Sec. 5.2.2. Secondly, an important perspective for future work is to focus on the probabilistic aspects of the model to improve its reliability in terms of probabilistic metrics as logScore or CRPS (Baggio and Muzy, 2024) and to compare it with other related approaches (for instance, non-parametric models based on normalizing flows (Marcille et al., 2024)). Additionally, our predictions could be improved by leveraging ensemble NWP forecasting, as pioneered by Rasp and Lerch (2018). This can potentially lead to better risk assessment for extreme events (Primo et al., 2024). Another promising direction involves incorporating additional data sources such as orography or surface roughness, that have been shown to influence the reliability of wind speed predictions (Rasp and Lerch, 2018; Veldkamp et al., 2021). The present study focuses on short-term forecasts (1–6 hours ahead) due to the use of the MeteoNet dataset. However, our model is inherently capable of handling longer forecast horizons so we plan to extend this approach to full intra-day (1-24 h) and day-ahead (>> 24 h) forecasts, by leveraging datasets that provide more frequent NWP updates over longer time ranges. Finally, advancements in deep learning architectures present promising opportunities for further refinement. While LSTMs and CNNs have demonstrated strong performance, recent advances in Transformer-based models have shown great potential for capturing long-range dependencies in spatiotemporal data (see, e.g., Ødegaard Bentsen et al. (2023)).
Appendix A The M-Rice probability distribution
The Rice probability distribution corresponds to the amplitude (i.e. the norm) of a two dimensional random vector which components are 2 independent Gaussian random variables of mean $\mu_{1}$ and $\mu_{2}$ and of same variance $\sigma^{2}$. The corresponding probability density function, $f_{R}(z,\nu,\sigma)$ is the following (Rice, 1945):
$f_{\text{R}}(y|\nu,\sigma)=\frac{y}{\sigma^{2}}e^{-\frac{y^{2}+\nu^{2}}{2\sigma^{2}}}\;I_{0}\Big(\frac{y\nu}{\sigma^{2}}\Big)$ (17)
where $\nu=\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}$ and $I_{0}(z)$ is the order zero modified Bessel function of the first kind.
Rice distribution has been extended by Baïle et al. (2011) to “Multifractal Rice” (M-Rice) distribution that accounts for the situation when, as observed in turbulence models, the variance $\sigma^{2}$ is itself stochastic with a log-normal distribution. The M-Rice distribution involves 3 parameters, namely the two Rice parameters coming from from Gaussian law $\nu$ and $\sigma^{2}$ and a supplementary parameter, denoted as $\lambda^{2}$ associated with the variance of the log-normal law. This parameter is referred to, in the literature on turbulence, as the “intermittency coefficient” (Frisch, 1995). The M-Rice probability density function (PDF) is then:
$f_{\text{MR}}(y|\nu,\sigma,\lambda^{2})=\frac{1}{\sqrt{2\pi\lambda^{2}}}\int e^{-\frac{\omega^{2}}{2\lambda^{2}}}f_{\text{R}}(y|\nu,\sigma e^{\omega})\;d\omega$
$=\!\frac{1}{\sqrt{2\pi\lambda^{2}}}\!\!\int\!\!\!\!e^{-\frac{\omega^{2}}{2\lambda^{2}}}\!\!\frac{y}{e^{2\omega}\sigma^{2}}e^{-\frac{y^{2}+\nu^{2}}{2e^{2\omega}\sigma^{2}}}I_{0}(\frac{y\nu}{e^{2\omega}\sigma^{2}})d\omega\;.$ (18)
@@ Appendix A The M-Rice probability distribution
As advocated in Baggio and Muzy (2024), this last formula is evaluated using the following Gauss-Hermite quadrature:
$\int_{-\infty}^{+\infty}e^{-y^{2}}f(y)\,dy\approx\sum_{i=1}^{n}w_{i}f(y_{i})$ (19)
@@
with $n=11$ that is sufficient for the purpose of this paper.
The M-Rice cumulative distribution function (CDF) function, useful to estimate the quantiles, can be obtained using the same quadrature formula and replacing the Rice PDF expression in (18) by the Rice CDF. This is also valid for the computation of the mean value: since for a Rice law of parameter $\nu$ and $\sigma^{2}$, the mean value is $\mu_{\text{R}}={\displaystyle\sigma{\sqrt{\frac{\pi}{2}}}\,\,L_{\frac{1}{2}}\left(-\frac{\nu^{2}}{2\sigma^{2}}\right)}$, where $L_{\frac{1}{2}}$ stands for the order $\frac{1}{2}$ Laguerre polynomial, the mean value of a M-Rice distribution reads:
$\mu_{\text{MR}}(\nu,\sigma,\lambda^{2})$ $=$ $\frac{\sigma}{\sqrt{2}}\int e^{-y^{2}}\,L_{\frac{1}{2}}\left(-\frac{e^{2\sqrt{2}\lambda y}\nu^{2}}{2\sigma^{2}}\right)\;dy$ (20)
$\simeq$ $\frac{\sigma}{\sqrt{2}}\sum_{i=1}^{n}w_{i}\,L_{\frac{1}{2}}\left(-\frac{e^{2\sqrt{2}\lambda y_{i}}\nu^{2}}{2\sigma^{2}}\right)\;.$
Open Research Section
Data Availability
The data used in this study originate from the MeteoNet database (Larvor et al., 2025), originally developed by Météo-France. A reference version used in this work is available on the Harvard Dataverse at https://doi.org/10.7910/DVN/NCKRZ2.
Code Availability
The source code used in this study, including data preprocessing functions, model implementations, analysis scripts, and a minimal example dataset, is publicly available under the MIT License at https://doi.org/10.5281/zenodo.15222910 (see Muzy and Baggio (2025)).
Acknowledgments
This work was supported by ANR grant SAPHIR project ANR-21-CE04-001403. The authors would like to acknowledge Marie-Laure Nivet (SPE) for insightful discussions and valuable assistance on matters related to open science.